OpenStudy (anonymous):

If we toss four fair coins, what is the probability that we get exactly four tails?

OpenStudy (anonymous):

A. 1/32 B. 1/8 C. 1/16 D. 1/20 :)

OpenStudy (anonymous):

(1/2)*(1/2)*(1/2)*(1/2) = (1/16)

OpenStudy (anonymous):

There are two possibilities We toss the coins 4 times So it would be 2*2*2*2=16 So you divide the probability by 1.